# Semi-classical imprints on quasinormal mode spectra

### v2

@@ Appendix C The radial wave equation
$\sigma(x)$ $=\frac{1}{(x-1)^{9}x^{2}}\left\{(x-1)^{4}(a_{1}r_{g}-1)\left[2r_{g}x^{5}\omega+(2r_{g}\omega-i)\right.\right.$
@@
$\left.+x(1-a_{1}r_{g})\left(a_{2}r_{g}^{2}+1\right)\left[1+4ir_{g}\omega\right.\right.$
@@
$\left.-(x-1)^{3}x\left(a_{2}r_{g}^{2}+1\right)\left[1+2ir_{g}\omega\right.\right.$
@@
$\tau(x)$ $=\frac{1}{(x-1)^{9}}\left\{(x-1)^{6}(x+1)\left(-a_{2}r_{g}^{2}-1\right)(1+2ir_{g}\omega)\right.$
@@
$\left.+\frac{(x-1)^{6}(a_{1}r_{g}-1)\left((x-1)x+2ir_{g}\left((x-1)^{2}x-1\right)\omega\right)}{x}\right.$
$\left.+(x-1)^{3}+(a_{1}r_{g}-1)\left(a_{2}r_{g}^{2}+1\right)\left[3x^{2}-1\right.\right.$
@@ Appendix D Accuracy of matrix method
Comparable accuracy to that which is demonstrated by (48) and (50) is obtained for higher-$l$ modes. We also compare the absolute and relative errors (in the static case) for the known values of the $l=0$ and $l=2$ modes. These values are obtained from the pseudo-spectral II algorithm with 40 polynomials, as reported in [59]. For the $l=0$ modes, $N>10$ provides a relative error better than $10^{-3}$, while a grid with $N>30$ is required to reach error below $10^{-5}$. The situation is improved for the (more relevant) $l=2$ mode, for which $N=27$ achieves an error as small as $10^{-10}$ relative to the pseudo-spectral II result.

### v3

@@ Appendix C The radial wave equation
$\sigma(x)$ $=\frac{1}{(x-1)^{9}x^{2}}\left\{(x-1)^{4}(\alpha_{1}-1)\left[2r_{g}x^{5}\omega+(2r_{g}\omega-i)\right.\right.$
@@
$\left.+x(1-\alpha_{1})\left(\alpha_{2}+1\right)\left[1+4ir_{g}\omega\right.\right.$
@@
$\left.-(x-1)^{3}x\left(\alpha_{2}+1\right)\left[1+2ir_{g}\omega\right.\right.$
@@
$\tau(x)$ $=\frac{1}{(x-1)^{9}}\left\{(x-1)^{6}(x+1)\left(-\alpha_{2}-1\right)(1+2ir_{g}\omega)\right.$
@@
$\left.+\frac{(x-1)^{6}(\alpha_{1}-1)\left((x-1)x+2ir_{g}\left((x-1)^{2}x-1\right)\omega\right)}{x}\right.$
$\left.+(x-1)^{3}+(\alpha_{1}-1)\left(\alpha_{2}+1\right)\left[3x^{2}-1\right.\right.$
@@ Appendix D Accuracy of matrix method
We provide some demonstrations of the accuracy and convergence of the matrix method towards known quasinormal frequencies. For example, with a uniformly spaced grid $x_{i}\in\{0,\tfrac{1}{4},\tfrac{2}{4},\tfrac{3}{4},1\}$, one obtains for the $l=m=0$ fundamental mode of the Schwarzschild black hole:
$\omega_{000}=0.21863-0.19826i$ (83)
This should be compared to the known frequency obtained through Leaver’s or pseudo-spectral methods:
$\omega_{000}=0.22091-0.20979i$ (84)
With instead a uniformly spaced grid $x_{i}\in\{0,\tfrac{1}{20},...,\tfrac{19}{20},1\}$, one obtains for the $l=m=0$ fundamental mode
$\omega_{000}=0.22088-0.20979i\ .$ (85)
Similar accuracy is obtained for higher-$l$ modes. We also compare the absolute and relative errors (in the static case) for the known values of the $l=0$ and $l=2$ modes. These values are obtained from the pseudo-spectral II algorithm with 40 polynomials, as in [59]. for the $l=0$ modes, $N>10$ provides a relative error better than $10^{-3}$, while a grid with $N>30$ is required to reach error below $10^{-5}$. The situation is improved for the (more relevant) $l=2$ mode, for which $N=27$ achieves an error as small as $10^{-10}$ relative to the pseudo-spectral II result.